# OrbitZoo: Multi-Agent Reinforcement Learning Environment for Orbital Dynamics

### v2

@@ 4 Novel Reinforcement Learning Challenges in OrbitZoo
In the emerging era of autonomous systems operating in the physical world, RL is evolving beyond synthetic benchmarks and addressing real-world constraints such as actuation limits, uncertainty, safety, and coordination. OrbitZoo is designed to support this shift. It provides a modular testbed to rigorously investigate RL challenges grounded in high-fidelity dynamics while abstracting just enough to enable algorithmic insight. Below, we outline key research questions that OrbitZoo helps explore.
@@
The environment supports decentralized, partially observable multi-agent setups via PettingZoo. These include tasks such as cooperative maneuvers, target chasing, and collision avoidance, scenarios rarely addressed in standard RL environments. OrbitZoo also enables centralized training and parallel simulations for scalability.
@@
Designing informative yet physically plausible reward signals is a key challenge in orbital control. OrbitZoo enables the use of parameterized reward functions, including orbital element targets (e.g., semi-major axis, inclination), fuel penalties, and safety-based margins, that enable flexible experimentation across shaped and sparse regimes. This supports research on credit assignment, reward misspecification, and curriculum learning in complex multi-agent environments.
@@
OrbitZoo’s modular setup supports adversarial scenarios (e.g., pursuit-evasion, intentional jamming), allowing RL agents to be evaluated in high-stakes conditions. Combined with partial observability and fuel-aware dynamics, this supports research in safe, robust policy design.
@@
OrbitZoo offers a real-time, integrated 3D visualization tool that aids policy inspection and failure case diagnosis. Although not a core contribution, this feature supports reproducibility and interpretability, and is summarized in Appendix A.
@@ 5 Experiments and Results
To evaluate the effectiveness of OrbitZoo in modeling diverse orbital missions and supporting RL methods, we conducted a series of experiments using several learning algorithms. These include the development of a mission analogous to one created in a different environment, orbital transfers (OTs), collision avoidance maneuvers (CAMs), a multi-agent geostationary orbit (GEO) constellation coordination problem using independent and federated learning, and a validation experiment comparing OrbitZoo’s simulations against real-world Starlink ephemeris data. In the following missions, we consider an RL agent as a satellite with limited maneuvering capabilities, learning a strategy (policy) to apply thrusts that maximize the expected return (Sec. 3.1).
@@ 5.2 Single-Agent Collision Avoidance Maneuver
CAM missions in RL are often focused on station-keeping while preventing collisions with other objects. Many existing approaches assume perfect knowledge of the current states of all bodies in the environment, simplifying the task to maximizing the Euclidean distance from other bodies while maintaining proximity to a nominal orbit. In reality, however, the states of these bodies are known with uncertainty, which must be accounted for in operational planning. In this mission, a more realistic approach to CAM is adopted by explicitly modeling state uncertainty and its evolution over time. While the environment handles propagation using high-fidelity dynamics that closely resemble real-world physics, the prediction of future states -— and the corresponding Probability of Collision (PoC) poc_akella -— relies on a simplified Newtonian attraction model, representing a low-fidelity simulation.
@@ Results.
In many existing CAM approaches, the RL agent is trained using discrete action spaces, since maneuvers are only required when a collision risk exists. Accordingly, a DQN was employed for training the agent. To investigate whether applying thrust in specific directions could lead to improved performance, the continuous version of PPO was also implemented. In this setup, the action space included a decision variable ($0\leq\delta\leq 1$) indicating whether thrust should be applied ($\delta>0.5$). For training, the environment used Newtonian attraction with atmospheric drag to model the actual dynamics. During evaluation, however, all available perturbations in OrbitZoo were enabled to assess the generalization capability of the trained agents. Results show that both algorithms learned effective policies under the (simplified) training dynamics, but PPO demonstrated superior performance under realistic conditions, exhibiting a greater ability to reduce the PoC.
@@ 5.5 Additional Experiments
For completeness, we present three additional experiments: (i) a mission implementation comparison with another environment (Appendix E.2), (ii) a chase-target scenario (Appendix E.4), where a satellite pursues a moving object in a higher orbit, and (iii) a single-agent collision avoidance task (Appendix E.5), where a satellite evades space debris in LEO. For each mission, we provide a brief overview, followed by the experimental setup, results, key challenges, and directions for future improvement.
@@ Conclusions
Through comparisons with existing environments and a diverse set of experiments—including a Hohmann transfer maneuver, a geostationary constellation coordination task, and validation using real-world Starlink ephemeris data—we demonstrated that RL policies trained in OrbitZoo can achieve near-optimal control while remaining consistent with realistic satellite behavior. These results establish OrbitZoo as a reliable benchmark for RL-based autonomy in space operations. To assess generalization and robustness, we employed multiple RL algorithms and different levels of realism, highlighting how continuous action spaces generally enhance performance under realistic orbital dynamics. Beyond RL, OrbitZoo serves as an open-source platform designed to support researchers in aerospace engineering, satellite operations, and machine learning. It provides a modular, high-fidelity environment for studying RL in realistic orbital settings. While grounded in orbital dynamics, its abstractions and challenge structures are representative of broader autonomy domains. The platform fosters reproducible experimentation and enables the exploration of autonomous decision-making for future space applications.
@@ A.2 Core Modules
(a) A system with 3 different Keplerian orbits.
(b) A system propagating several Starlink satellites.
(c) A single-agent mission: Follow Leader.
(d) A multi-agent mission: Constellation in GEO.
Figure 5: Frames of different systems on OrbitZoo’s interface.
@@ Appendix E Experiments
To evaluate OrbitZoo’s capabilities, we designed a series of missions utilizing algorithms ranging from foundational to state-of-the-art (as discussed in Section E.1). The first mission (Section E.2) compares the OrbitZoo environment to D. Kolosa’s implementation for a general orbit change maneuver using DDPG, highlighting how the tools and flexibility provided by OrbitZoo can support the development of both existing and novel, more complex missions – without the technicalities typically present in orbital dynamics.
Next, two missions (Sections E.3 and E.4) focus on orbital transfers (OTs) using PPO, where agents learn strategies to reach static and dynamic targets, respectively, while considering real-world factors such as perturbative forces and mass loss due to fuel consumption.
@@ E.1.3 DDQN
E.2 OrbitZoo vs. SOTA: Environment Comparison
@@ E.2.1 Definition
Many RL missions and environments developed by other authors (as shown in Table 1) are not publicly available, hindering reproducibility and making it challenging to assess the complexity of their implementation. Among the few exceptions is the work by D. Kolosa [40] and A. Herrera [31], whose code is openly accessible and supports the results presented. In this context, we replicate Kolosa’s first mission to demonstrate how OrbitZoo can be configured to achieve similar objectives—more easily and with greater flexibility. A reproduction of Herrera’s mission is included in the supplementary code.
@@ E.2.4 Challenges and Future Improvements
E.3 Hohmann Maneuver
@@ E.3.3 Results
E.3.4 Challenges and Future Improvements
@@
E.4 Chase Target
E.4.1 Definition
@@ E.4.3 Results
E.4.4 Challenges and Future Improvements
@@
E.5 Collision Avoidance
E.5.1 Definition
@@ 6 $(0,0,0,0)$ Nothing
E.5.4 Challenges and Future Improvements
@@
E.6 GEO Constellation
E.6.1 Definition
Unlike Low Earth Orbit (LEO) and Medium Earth Orbit (MEO), which occupy defined altitude ranges, the Geostationary Earth Orbit (GEO) is a specific orbit located 35786 km above Earth’s equator. This orbit is significant because any object within it maintains an orbital period equal to Earth’s rotational period, making it appear stationary to observers on Earth.
@@ E.6.3 Results
E.6.4 Challenges and Future Improvements
@@ F.2 Residual Scenarios
Figures 30 to 33 illustrate the residuals for these scenarios. The overlapping region provides the most reliable assessment of the propagator’s accuracy, as it is based on the most recent and accurate ephemeris data.

### v3

# OrbitZoo: Real Orbital Systems Challenges for Reinforcement Learning

@@ 4 Novel Reinforcement Learning Challenges in OrbitZoo
In the emerging era of autonomous systems operating in the physical world, RL is evolving beyond synthetic benchmarks and addressing real-world constraints such as actuation limits, uncertainty, safety, and coordination. OrbitZoo is designed to support this shift. It provides a modular testbed to rigorously investigate RL challenges grounded in high-fidelity dynamics while abstracting just enough to enable algorithmic insight. Tab. 2 summarizes the main RL challenges addressed by OrbitZoo, which are discussed in detail below.
Table 2: Summary of reinforcement learning challenges and how OrbitZoo addresses them.
Research Gap Limitations of Standard Environments Enabled by OrbitZoo
Sim-to-Real Transfer Neglects real-world orbital perturbations and environmental uncertainties. Realistic dynamics with harmonic gravity fields, drag, SRP, third body forces, uncertainties and variable-step integration methods.
Continuous Control Impulsive or unconstrained control models. Time-continuous thrust with fuel and actuation constraints.
Multi-Agent Coordination Mostly single-agent, fully observable setups. Decentralized agents with partial observability via PettingZoo.
Reward Design Dense, artificial reward signals. Flexible, physically grounded reward functions.
Safety and Adversarial Learning Rarely models safety or adversarial dynamics. Supports pursuit–evasion and safe maneuvering tasks.
Visualization Limited (2D) or no visualization for debugging. Real-time 3D inspection of agent behavior.
@@
Real-world orbital operations often involve multiple autonomous spacecraft interacting under limited information – for example, in formation flying, debris tracking, or pursuit/evasion scenarios. Each agent must make decisions based only on local or delayed observations, while the overall behavior emerges from their collective interaction. OrbitZoo enables such decentralized and partially observable setups through its PettingZoo-based architecture, where each agent receives its own observation and action space while sharing the same dynamic environment.
This design allows investigation of both cooperative and competitive MARL paradigms. Cooperative tasks include coordinated thrusting for formation maintenance or rendezvous, while adversarial ones involve pursuit, evasion, or interference. OrbitZoo also supports centralized training with decentralized execution (CTDE), where a global critic can access full system information during learning, but each agent acts autonomously at runtime. Combined with its support for parallel rollouts, this makes OrbitZoo a scalable testbed for studying emergent coordination, communication strategies, and robustness in multi-agent orbital control – domains that remain largely unexplored in standard RL benchmarks.
@@
Designing informative yet physically grounded reward signals is a central challenge in orbital control. Rewards must capture long-term mission objectives while remaining interpretable within physical constraints. OrbitZoo provides a flexible reward framework that integrates both inter-body and body-specific metrics, enabling the formulation of rich and diverse learning objectives. Inter-body quantities – such as relative distance or line-of-sight conditions – allow the definition of cooperative or competitive multi-agent tasks. In parallel, each body exposes individual physical attributes such as fuel consumption and mass variation, which can be incorporated into agent-specific reward functions.
This modular reward interface enables experimentation across dense and sparse regimes, as well as multi-objective trade-offs between performance, safety, and efficiency. It supports studies on credit assignment, reward misspecification, and curriculum learning curriculum-learning in complex multi-agent orbital environments, where delayed effects and coupled dynamics make the reward landscape inherently nontrivial.
@@
OrbitZoo’s modular setup supports adversarial and safety-critical scenarios, such as pursuit–evasion, intentional jamming, or collision avoidance. The environment includes inter-body metrics like the probability of collision (POC) and body-specific measures of propagation uncertainty, allowing agents to quantify and respond to risk in real time. Combined with partial observability and fuel-aware dynamics, these features enable research on robust policy design, safe exploration, and adversarial resilience in continuous orbital domains.
@@
OrbitZoo offers a real-time, integrated 3D visualization tool that aids policy inspection and failure case diagnosis. Although not a core contribution, this represents, to the best of our knowledge, the first Python-based RL framework with a built-in, real-time orbital visualization interface, as shown in Fig. 1 and summarized in Appendix A. This feature enhances interpretability, reproducibility, and debugging by providing an intuitive link between learned behaviors and the underlying orbital dynamics.
(a) A system with 3 different Keplerian orbits.
(b) A system propagating several Starlink satellites.
(c) A single-agent mission: Follow Leader.
(d) A multi-agent mission: Constellation in GEO.
Figure 1: Frames of different systems on OrbitZoo’s interface.
@@ 5 Experiments and Results
To evaluate the effectiveness of OrbitZoo in modeling diverse orbital missions and supporting RL methods, we conducted a series of experiments using several learning algorithms. These include the development of a mission analogous to one created in a different environment, orbital transfers (OTs), collision avoidance maneuvers (CAMs), a multi-agent geostationary orbit (GEO) constellation coordination problem using independent and federated learning, and a validation experiment comparing OrbitZoo’s simulations against real-world Starlink ephemeris data. In the following missions, we consider an RL agent as a satellite with limited maneuvering capabilities, learning a strategy (policy) to apply thrusts that maximize the expected return (Sec. 3.1). Details can be found in Appendix E.
@@ 5.2 Single-Agent Collision Avoidance Maneuver
CAM missions in RL are often focused on station-keeping while preventing collisions with other objects. Many existing approaches assume perfect knowledge of the current states of all bodies in the environment, simplifying the task to maximizing the Euclidean distance from other bodies while maintaining proximity to a nominal orbit. In reality, however, the states of these bodies are known with uncertainty, which must be accounted for in operational planning. In this mission, a more realistic approach to CAM is adopted by explicitly modeling state uncertainty and its evolution over time. While the environment handles propagation using high-fidelity dynamics that closely resemble real-world physics, the prediction of future states – and the corresponding Probability of Collision (PoC) poc_akella – relies on a simplified Newtonian attraction model, representing a low-fidelity simulation.
@@ Results.
In many existing CAM approaches, the RL agent is trained using discrete action spaces, since maneuvers are only required when a collision risk exists. Accordingly, a DQN was employed for training the agent. To investigate whether applying thrust in specific directions could lead to improved performance, the continuous version of PPO was also implemented. In this setup, the action space included a decision variable ($0\leq\delta\leq 1$) indicating whether thrust should be applied ($\delta>0.5$). For training, the environment used Newtonian attraction with atmospheric drag to model the actual dynamics. During evaluation, however, all available perturbations in OrbitZoo were enabled to assess the generalization capability of the trained agents. Results show that both algorithms learned effective policies under the (simplified) training dynamics, but PPO demonstrated superior performance under realistic conditions, exhibiting a greater ability to reduce the PoC. More details are presented in Appendix E.6.
@@ 5.5 Additional Experiments
Many RL missions and environments developed by other authors (as shown in Table 1) are not publicly available, hindering reproducibility and making it challenging to assess the complexity of their implementation. Among the few exceptions is the work by D. Kolosa Kolosa2019 and A. Herrera herrera2020 , whose code is openly accessible and supports the results presented. In this context, we replicate both authors first missions in Appendix E.2 and Appendix E.3, respectively, to illustrate how OrbitZoo provides a flexible framework for developing and analyzing such missions. For completeness, we present an additional experiment: a chase-target scenario (Appendix E.5), where a satellite pursues a moving object in a higher orbit. For each mission, we provide a brief overview, followed by the experimental setup, results, key challenges, and directions for future improvement.
@@ Conclusions
Through comparisons with existing environments and a diverse set of experiments – including a Hohmann transfer maneuver, a geostationary constellation coordination task, and validation using real-world Starlink ephemeris data – we demonstrated that RL policies trained in OrbitZoo can achieve near-optimal control while remaining consistent with realistic satellite behavior. These results establish OrbitZoo as a reliable benchmark for RL-based autonomy in space operations. To assess generalization and robustness, we employed multiple RL algorithms and different levels of realism, highlighting how continuous action spaces generally enhance performance under realistic orbital dynamics. Beyond RL, OrbitZoo serves as an open-source platform designed to support researchers in aerospace engineering, satellite operations, and machine learning. It provides a modular, high-fidelity environment for studying RL in realistic orbital settings. While grounded in orbital dynamics, its abstractions and challenge structures are representative of broader autonomy domains. The platform fosters reproducible experimentation and enables the exploration of autonomous decision-making for future space applications.
@@ Appendix E Experiments
To evaluate OrbitZoo’s capabilities, we designed a series of missions utilizing algorithms ranging from foundational to state-of-the-art (as discussed in Section E.1). The first mission (Section E.2) compares the OrbitZoo environment to D. Kolosa’s implementation for a general orbit change maneuver using DDPG, highlighting how the tools and flexibility provided by OrbitZoo can support the development of both existing and novel, more complex missions – without the technicalities typically present in orbital dynamics. Then, Herrera’s [31] propagation method is compared to the one used in OrbitZoo, which is used to develop a station-keeping mission in LEO (Section E.3).
Next, two missions (Sections E.4 and E.5) focus on orbital transfers (OTs) using PPO, where agents learn strategies to reach static and dynamic targets, respectively, while considering real-world factors such as perturbative forces and mass loss due to fuel consumption.
@@ E.1.3 DDQN
E.2 OrbitZoo vs. SOTA: Kolosa Comparison
@@ E.2.1 Definition
Kolosa [40] developed a mission in Medium Earth Orbit (MEO) where a spacecraft learns to perform a general orbit change maneuver using DDPG. In this section, we replicate Kolosa’s mission within OrbitZoo.
@@ E.2.4 Challenges and Future Improvements
E.3 OrbitZoo vs. SOTA: Herrera Comparison
To compare OrbitZoo’s capabilities with those of existing environments that offer publicly available code, we extend our analysis by implementing the mission proposed by Herrera in [31], consisting of a station-keeping mission in LEO.
E.3.1 Definition
Herrera developed a reinforcement learning environment for station-keeping in Low Earth Orbit (LEO). While Keplerian orbits describe a well-defined elliptical path around a central body, real-world satellites are influenced by perturbative forces beyond gravity, causing them to drift from their nominal orbits over time. Station-keeping missions aim to apply strategic control – through attitude adjustments or thrust maneuvers – to maximize the duration a satellite stays within its designated orbit. This task becomes especially challenging and critical in LEO due to the significant impact of atmospheric drag.
E.3.2 Environment Setup
Objective and Environment Characteristics.
In this mission, Herrera employed PPO to train a satellite to control its thrust for station-keeping. The environment’s dynamics were manually implemented using a 4th-order Yoshida integrator, which computes the satellite’s next position and velocity based on gravitational and drag forces. The satellite began on a circular orbit at an altitude of 550 km above Earth’s surface, with the objective of maintaining that altitude for as long as possible without deviating by more than 1 meter.
Herrera also clearly defined the episode termination conditions: (1) each episode has a maximum duration of 800 steps, with each step representing 1 second; (2) if the satellite does not perform any maneuvers, it exceeds the threshold after approximately 200 steps due to natural drift; and (3) the satellite is provided with enough fuel to apply maximum thrust for up to 125 steps. For a mission requiring this level of precision, the differences in dynamic modeling are critical. As stated by Herrera: "The 4th Order Yoshida integrator is used as the default integrator as there is a balance of performance and error. RK4, while generally having less error, is not a semantic integrator and requires the calculation of the acceleration 4 times while the Yoshida integrator only requires 3.". Given the difference in integration methods – Herrera employing the faster but less precise Yoshida integrator, and OrbitZoo using the more accurate Dormand-Prince integrator (a RK4 integrator with adaptive time steps) – we conduct a direct comparison between the values produced by each environment over the course of a full episode, starting from the same initial position and velocity. As shown in Figure 15, the difference in position after approximately 13 minutes is around 25 meters, while the difference in velocity is around 0.08 meters per second. With this comparison, we note that the one meter threshold initially defined for episode termination is extremely strict under realistic dynamic conditions. The code needed to run this comparison is provided in the supplementary material.
(a) Euclidean distance between Herrera’s and OrbitZoo’s propagated position, per second.
(b) Euclidean distance between Herrera’s and OrbitZoo’s propagated velocity, per second.
Figure 15: Comparison between Herrera’s and OrbitZoo’s natural propagation: position and velocity differences over time.
Following Herrera’s setup, we designed a similar mission in which the agent has an initial mass of 100 kg, including 75 kg of fuel. The satellite is modeled as a perfect sphere with a radius of 16.8 meters, a drag coefficient of 2.123 and is equipped with a thruster characterized by a specific impulse of $I_{\text{sp}}=0.0067$ s. While this configuration is not physically realistic, it was chosen to maintain consistency with Herrera’s termination conditions (2) and (3). Notably, in OrbitZoo, atmospheric drag is not solely determined by the satellite’s current state (as in Herrera’s implementation) but also incorporates historical data, allowing drag to vary dynamically based on the current moment.
Action Space.
To simplify the problem, Herrera limits the action space to the orbital plane, and considers a polar thrust representation $(T,\theta)$. However, the action the agent performs is not directly the thrust being applied, but the variation applied to the current thrust, creating a smoother and more realistic maneuver: $a_{t}=(\Delta T,\Delta\theta)$. While the maximum thrust is characterized by $(T_{\text{max}},\theta_{\text{max}})=(0.04,2\pi)$, the action space is limited to a fraction of that change: $(\Delta T_{\text{max}},\Delta\theta_{\text{max}})=(T_{\text{max}}/50,\theta_{\text{max}}/6)$.
Observation Space.
Similarly to Herrera, we use Cartesian coordinates to represent the current position $r\in\mathbb{R}^{2}$ and velocity $\dot{r}\in\mathbb{R}^{2}$ of the satellite. Additionally, we also define the distance to the nominal altitude, $r_{\text{target}}=|\|r\|-r_{\text{nominal}}|$ and nominal velocity, $\dot{r}_{\text{target}}=|\|\dot{r}\|-\dot{r}_{\text{nominal}}|$, where
$r_{\text{nominal}}=550\text{ km}\quad\quad\dot{r}_{\text{nominal}}=\sqrt{\frac{\mu_{E}}{r_{\text{nominal}}}},$
with $\mu_{E}$ representing the gravitational parameter of Earth. The observation $o_{t}\in\mathbb{R}^{8}$ consists of: $o_{t}=(r/r_{\text{nominal}},\dot{r}/\dot{r}_{\text{nominal}},r_{\text{target}},\dot{r}_{\text{target}},\theta,T)$.
Reward Function.
In order to encourage the agent to both not run out of fuel and stay within the orbital termination threshold, the reward function is similar to Herrera’s:
[…]
where $f$ is the current available fuel in kg, and $t$ is the current step within the episode.
E.3.3 Results
Similarly to Herrera, we employed PPO to trained the agent, with the algorithm hyperparameters being shown in Table 9. If the agent can consistently exceed the baseline of 200 steps – representing the duration a satellite remains within tolerance without maneuvers – it demonstrates successful station-keeping. Figure 16 illustrates that, despite the narrow 1-meter tolerance from the nominal orbit (together with the more realistic perturbations), the agent effectively learned to maintain station-keeping by surpassing those 200 steps consistently.
Table 9: OrbitZoo vs. Herrera: Training hyperparameters.
Parameter Value
Actor learning rate 0.0001
Critic learning rate 0.001
GAE $\lambda$ 0.95
Epochs 5
Discount factor ($\gamma$) 0.99
Clip ($\epsilon$) 0.03
Initial Standard Deviation 0.5
Standard Deviation Decay Rate 10000 steps
Standard Deviation Decay Amount 0.05
Minimum Standard Deviation 0.05
Experiences for Training 800
Batch Size 64
Figure 16: Steps per episode, averaged by a sliding window of 10 episodes. If no maneuvers are executed, the satellite exits the allowed tolerance range by step 200.
E.3.4 Challenges and Future Improvements
The agent did not achieve the same level of performance as in Herrera’s environment, primarily due to differences in the dynamics, as demonstrated earlier. Additionally, the PPO implementations differed – one using TensorFlow and the other PyTorch – with variations in hyperparameters.
This attempt to replicate Herrera’s environment suggests that OrbitZoo could benefit from using faster propagators. While these may be less precise, they are suitable for RL missions where extreme precision is not critical. As Herrera points out, continuous but deterministic algorithms (such as DDPG and TD3) may be better suited for this mission due to the precision required in controlling thrust maneuvers – something that is challenging to achieve with PPO.
E.4 Hohmann Maneuver
@@ E.4.3 Results
E.4.4 Challenges and Future Improvements
@@
E.5 Chase Target
E.5.1 Definition
@@ E.5.3 Results
E.5.4 Challenges and Future Improvements
@@
E.6 Collision Avoidance
E.6.1 Definition
@@ 6 $(0,0,0,0)$ Nothing
E.6.4 Challenges and Future Improvements
@@
E.7 GEO Constellation
E.7.1 Definition
Unlike LEO and MEO, which occupy defined altitude ranges, the Geostationary Earth Orbit (GEO) is a specific orbit located 35786 km above Earth’s equator. This orbit is significant because any object within it maintains an orbital period equal to Earth’s rotational period, making it appear stationary to observers on Earth.
@@ E.7.3 Results
E.7.4 Challenges and Future Improvements
@@ F.2 Residual Scenarios
Figures 32 to 35 illustrate the residuals for these scenarios. The overlapping region provides the most reliable assessment of the propagator’s accuracy, as it is based on the most recent and accurate ephemeris data.